OpenStudy (anonymous):

A 0.016402-mol sample of an organic compound was burned in oxygen in a bomb calorimeter. The temperature of the calorimeter increased from 26.3oC to 27.3oC. If the heat capacity of the calorimeter is 4.28 kJ (oC)-1, then what is the constant volume heat of combustion of this compound, in kilojoules per mole?

OpenStudy (aaronq):

i'm not sure if the heat capacity given is in moles or grams, you have to double check that: use: \(q=n*C_p*\Delta T\) or \(q=m*C_p*\Delta T\) , depending on the units of \(C_p\). The scale your result to the amount of energy per 1 mole.